OpenStudy (anonymous):

Need step-by-step assistance for: 4w^2=-20w-25

OpenStudy (anonymous):

I know to convert it so ax^2+bx+c=0

OpenStudy (anonymous):

4w^2+20w+25=0

OpenStudy (anonymous):

convert it to*

OpenStudy (gorv):

4w2 +20w+25=0

OpenStudy (gorv):

4w2+10w+10w+25=0

OpenStudy (anonymous):

Okay, that's what I was having troules with, strene. I don't understand that- the 20w is a negative (-20w) so how can you make that automatically a positive?

OpenStudy (anonymous):

4w^2=2w.2w 25=5.5 (2w+5)(2w+5)=0

OpenStudy (gorv):

2w(2w+5)+5(2w+5)=0 (2w+5)(2w+5)=0

OpenStudy (anonymous):

please delete the further steps..... i need to follow through it step by step..

OpenStudy (gorv):

u can see it is a perfect square

OpenStudy (gorv):

just take -20w 0n other side

OpenStudy (anonymous):

if you pass a variable to the other side of equation, it changes sign - becomes +, + becomes -

OpenStudy (anonymous):

right, i know that. i thought there was something else to it